Show me how to slove the equation: 5/2t - t = 3 + 3/2t
Answer by edjones(8007)

5/2t - t = 3 + 3/2t
3/2t=3+3/2t
subtract 3/2t from each side: 0=3
the equation has no solution.
Ed